OpenStudy (anonymous):

PLEASE HELP WILL MEDEL AND FAN Which set is closed under multiplication? A. {0, 1, 2} B. {1, 2, 3} C. None of the sets are closed under multiplication. D. {1, 2}

OpenStudy (amistre64):

what does it mean for a set to e closed under an operation?

OpenStudy (albert0898):

Closed under multiplication means that any member of the set times any other member of the set (including itself) yields another member of the set. A: 0 * 0 = 0 , 0 * 1 = 0, 1 * 1 = 1, 2 * 2 = 4, 2 * 1 = 2 NOPE B: Now apply the same concept to the other sets, if you find on combination that yields something outside the set, then it is no closed. C: D: Now apply the same concept to the other sets, if you find on combination that yields something outside the set, then it is no closed.

OpenStudy (albert0898):

C is the right answer.

OpenStudy (amistre64):

since this site is not an answering service, please tutor and involve the asker in the solution process as opposed to simply doing all the work/thinking for them.

OpenStudy (albert0898):

@worker Basically, when multiplying the terms together and with itself, you can't get a result outside of the given set. @amistre64 I apologize. I did provide insight though

OpenStudy (amistre64):

you did yes :) i like that you are capable of giving a cogent answer.